X must borrow $2000 for 2 years. He is offered the money at: (A) 5% compounded quarterly (B) % 8 3 5 compounded annually (C) % 2 1 5 simple interest Which offer should he accept


1
Expert's answer
2021-11-16T09:27:48-0500

A)

"FV=2000(1+0.05/4)^{2\cdot4}=\$2208.97"


B)

"FV=2000(1+0.0835)^{2}=\$2347.94"


C)

"FV=2000(1+0.0215\cdot2)=\$2086"


Since "FV=\$2086" is minimal, he should accept offer (C).